OpenStudy (anonymous):

Find the value of the expression. [sin(70 degrees)/cos(-430 degrees)]+tan(-70degrees)

OpenStudy (anonymous):

@butterflydreamer Hey sorry, do you mind helping me with a couple more, I REALLY like your method of teaching.

OpenStudy (jhannybean):

well, negative angles means you're counting counter clockwise.

OpenStudy (jhannybean):

So you'd want to find the coterminal angle of -430 by subtracting \(-360^\circ\) till you find anangle between \(0~,~ 2\pi\) I believe.

OpenStudy (jhannybean):

That's one method. You can also find the radian measure by \[-430^\circ \cdot \frac{\pi}{180^\circ}\]